# This draft: December 17, 2025 First draft: March 5, 2025

### v1

@@ 7 Conclusion
A Consensus-Bottleneck Asset Pricing Model”  [^1] This paper is a revised version of Master’s thesis by Changeun Kim titled “A Consensus-Bottleneck Asset Pricing Model”, submitted to the Department of Industrial and Management Engineering, POSTECH, Korea. We would like to thank Hyeng Keun Koo, Kwangmin Jung, Dojoon Park (discussant), JinGi Ha, Jeonggyu Huh, Kyoung-Kuk Kim (discussant), Thummim Cho, and seminar participants at the 2024 Spring Joint Conference of Korean Operations Research and Management Science Society and Korean Institute of Industrial Engineers, 2025 Asia-Pacific Association of Finance International Conference, 2025 Korean Finance Association Fall Conference, 2025 4th Workshop on Financial Mathematics and Engineering (Pusan National University), 2025 Korea Derivatives Association Fall Conference, 2025 Annual Conference on Asia-Pacific Financial Markets (CAFM), for helpful discussions and insightful comments. This work was supported by the Ministry of Education of the Republic of Korea and the National Research Foundation of Korea (NRF-2023R1A2C2003927). Jang (E-mail: bonggyujang@postech.ac.kr) is at Department of Industrial and Management Engineering, POSTECH, Korea University Business School; Jeong (E-mail: younwoo48@postech.ac.kr) is at Graduate School of Artificial Intelligence, POSTECH; Kim (E-mail: changeun120@postech.ac.kr) is at Department of Industrial and Management Engineering, POSTECH. Correspondence concerning this article should be addressed to Changeun Kim, Department of Industrial and Management Engineering, POSTECH, Pohang 37673, Republic of Korea. E-mail: changeun120@postech.ac.kr.
@@
The Internet Appendix is organized as follows. Section A discusses data preprocessing procedures applied prior to estimation and evaluation. Section B provides details on the implementation and architectural choices of the neural network used throughout this study. Section C reports additional empirical results for robustness and supplementary insights. Section D presents the list of variables comprising the dataset used in this study.
Appendix A Data Preprocessing
@@ A.3 Data imputation
However, for variables that represent growth rates or changes in firm characteristics,  [^16] Variables marked with asterisks in Table LABEL:tab:firm-level-predictors of Internet Appendix D we apply time-series mean imputation. This method captures the inherent continuity and trend in firm-specific dynamics, producing more realistic estimates than cross-sectional averaging. Relying repeatedly on the last observed value for growth-related factors could falsely imply persistence or monotonic trends, misrepresenting the inherently dynamic nature of such variables.
@@ A.4 Data normalization
Appendix B Implementing Neural Network for Asset Pricing
B.1 Activation functions
Rectified Linear Unit (ReLU) is frequently chosen in various machine learning applications due to its computational simplicity and efficiency, as also evidenced by its usage in empirical asset pricing research employing deep learning architectures such as gu2020empirical and chen2024deep. In contrast to activations such as SoftMax or Sigmoid functions, ReLU has demonstrated comparable performance while offering faster computational speed. Moreover, ReLU is particularly valued for its ability to address the gradient vanishing problem by consistently producing non-negative gradients for positive inputs. The Figure B.1 illustrates a graphical form of ReLU function.
Figure B.1: Rectified Linear Unit (ReLU) activation function.
@@ B.1 Activation functions
B.2 Stabilized learning
Stabilizing neural network training is essential for robust estimation, particularly when deploying deep learning models in high-stakes decision-making contexts. The following sections detail the regularization methods and machine learning techniques we employ to enhance the stability of the learning process. [^17] Detailed parameter settings are provided in Table LABEL:tab:hyper-setting. Despite that this topic is related to experimental factors rather than based on theoretical backgrounds, it is still worthy to discuss since that not only providing these information is crucial for reproducing the results, but also ensures a clear understanding of deep learning techniques. This is particularly important in fields like financial economics, where such concepts may not yet be widely understood or adopted. As such, this discussion is not only relevant but vital for integrating advanced computational methods into financial economic research effectively.
@@ B.2 Stabilized learning
Adaptive Moment Estimation (ADAM) is an optimization algorithm proposed by kingma2017adam, which is commonly used in practice of training deep learning models. It combines ideas from both momentum optimization and RMSprop (hinton2012neural), making it well-suited for optimizing non-convex objective functions commonly encountered in neural network training. ADAM maintain the first moment and the second moment of the gradients as two separate moving averages. These moving averages are used to adaptively update the parameters of the model during training. ADAM automatically adjusts the learning rate for each parameter based on the magnitude of the gradients and the accumulated past gradients, allowing it to converge quickly and efficiently in practice.
Learning rate scheduling is a technique used to dynamically adjust the learning rate during training to improve optimization performance. Instead of using a fixed learning rate throughout the training process, learning rate scheduling gradually decrease the learning rate over time, allowing the model to fine-tune its parameters more effectively as training progresses. By annealing the learning rate, learning rate scheduling helps to prevent the optimization process from getting stuck in local minima, which happens surprisingly often as discussed in the next section. Although ADAM inherently adjusts the learning rate, combining it with learning rate scheduling further optimizes performance by refining the initial learning rate. For training CB-APM, we use the ReduceLROnPlateau scheduler provided in PyTorch (See the official document provided by PyTorch for more details ReduceLROnPlateau), which monitors validation performance and reduces the learning rate by a specified factor.
@@ B.3 Model hyperparameter
Given these considerations, designing an appropriate hyperparameter optimization problem for asset pricing is important, which we leave it as an attractive and also challenging future research topic. In this paper, we rather summarize the hyperparameter “choice” that produces reasonable results as practical guidelines for finding a rational setting. The list of all the hyperparameters under considerations are provided in Table LABEL:tab:hyper-setting.
Table B.1: Hyperparameter settings for CB-APM and Autoencoder.
@@
Table B.1: Hyperparameter settings for CB-APM and Autoencoder (cont’d).
@@
Appendix C Additional Results
C.1 Forecast horizon analysis
Table C.1: Out-of-Sample $R^{2}$ for Stock Return and Consensus Approximations.
@@ C.1 Forecast horizon analysis
Table C.2: Out-of-Sample $R^{2}$ for Stock Return and Consensus Approximations.
@@
Table C.3: Out-of-Sample $R^{2}$ for Stock Return and Consensus Approximations.
@@
Figure C.1: Out-of-sample $R^{2}$ of return predictions and consensus approximations.
Panel (a), (b), and (c) presents results for the monthly, quarterly, and semiannual returns respectively.
Figure C.2: Out-of-sample $R^{2}$ by testing period under expanding window evaluation.
This figure reports monthly $R^{2}$ of monthly stock return and consensus prediction by period for naïve neural network ($\lambda=0$, left) and the best-performing model ($\lambda=0.03$, right.
Figure C.3: Out-of-sample $R^{2}$ by testing period under expanding window evaluation.
This figure reports monthly $R^{2}$ of quarterly stock return and consensus prediction by period for naïve neural network ($\lambda=0$, left) and the best-performing model ($\lambda=0.65$, right.
Figure C.4: Out-of-sample $R^{2}$ by testing period under expanding window evaluation.
This figure reports monthly $R^{2}$ of semiannual stock return and consensus prediction by period for naïve neural network ($\lambda=0$, left) and the best-performing model ($\lambda=0.18$, right.
@@
Tables C.1C.3 and Figures C.1C.4 present the out-of-sample $R^{2}$ results and their period-by-period decomposition for these horizons. For shorter horizons such as one month, return predictability remains marginal, with $R^{2}$ values close to zero and even slightly negative at higher values of $\lambda$, consistent with the well-documented difficulty of forecasting near-term returns. In this regime, increasing $\lambda$ intensifies the interpretability-accuracy trade-off: while consensus approximation improves monotonically, return $R^{2}$ declines, suggesting that allocating more weight to consensus modeling diverts representational capacity away from short-horizon return-specific signals.
@@
C.2 Properties of the joint optimization
@@
Figure C.5 illustrates the in-sample MSE dynamics of the CB-APM under varying values of the hyperparameter $\lambda$, separately for monthly and annual forecasting horizons. Each panel presents two curves: the left axis depicts the in-sample MSE of stock return predictions, while the right axis reports the average in-sample MSE for consensus variable approximation. The first panel corresponds to the monthly horizon (Figure 5(a)), and the second panel presents the results for the annual horizon (Figure 5(b)).
@@
Figure C.5: In-sample MSE of return and consensus approximations.
This figure plots the in-sample mean squared error (MSE) of stock return (left) and consensus approximation (right) for different $\lambda$ settings. Panel (a) reports results for the monthly return, and Panel (b) for the annual return. These plots illustrate how forecasting horizons and $\lambda$ values govern the trade-off between predictive accuracy and consensus reconstruction in the joint loss function.
The observed patterns reveal a striking divergence between the two horizons. For monthly returns, the stock return MSE exhibits a U-shaped trajectory: it initially decreases slightly for small values of $\lambda$ but increases steadily thereafter, suggesting a trade-off between return prediction accuracy and consensus approximation performance. This pattern aligns with the theoretical role of $\lambda$ in the loss function in equation (4), which explicitly prioritizes consensus approximation as its value increases. Placing greater weight on $L_{C}$ directs more representational capacity of the network toward modeling analyst consensus at the expense of direct return prediction. This is consistent with the interpretability-accuracy trade-off widely documented in the interpretable machine learning literature (e.g., rudin2019stop), where models constrained to capture auxiliary structure or explanatory variables tend to sacrifice marginal predictive performance in favor of enhanced interpretability or alignment with economic reasoning.
@@
C.3 Structure of the learned macroeconomic representations
@@
Figure C.6 illustrates the two-dimensional principal component projection of the 32-dimensional latent state vectors produced by the macroeconomic autoencoder, color-coded by month and annotated with January observations for selected years. This visualization highlights how the autoencoder successfully encodes macroeconomic conditions into a smooth, low-dimensional manifold that evolves coherently over time. The trajectory of the latent vectors follows a clear temporal progression, demonstrating that the learned embedding captures the gradual transitions and structural shifts in the U.S. macroeconomic environment across the sample period.
Figure C.6: PCA projection of autoencoder latent state variables.
@@
A notable feature of this representation is its ability to distinguish major economic regimes. The grey dashed ovals in Figure C.6 correspond to periods classified as recessions by the National Bureau of Economic Research (NBER); the early 2000s recession (2001Q1–Q4) and the Global Financial Crisis (2007Q4–2009Q2). During these intervals, the latent vectors exhibit marked departures from their preceding trajectories, forming clusters that are distinct from surrounding expansionary phases. This pattern indicates that the autoencoder embedding effectively internalizes macroeconomic shocks and regime shifts, producing representations that align with well-established business cycle chronologies without direct supervision from recession labels.
@@
Recent deep factor models make it clear that neural networks can extract a parsimonious set of latent factors from high‐dimensional financial and macroeconomic data; these latent variables then drive improved asset pricing and predictive performance (see, for example, feng2018deep; gu2021autoencoder; chen2024deep). Our macroeconomic autoencoder performs a closely related function for aggregate time‐series data: it distills hundreds of macro indicators into a smooth latent trajectory that aligns with well‐known business‐cycle chronologies and regime shifts. While many prior studies emphasize quantitative performance metrics and offer only limited visual exploration of their latent factors, the clear temporal patterns in Figure C.6 demonstrate that CB‐APM’s autoencoder uncovers economically meaningful state dynamics from complex data.
@@
C.4 Further robustness checks
C.4.1 Sensitivity of autoencoder performance to latent dimensionality
@@
Table C.4: Out-of-Sample $R^{2}$ for Stock Return and Consensus Approximations for Different Autoencoder Dimensions.
@@
C.4.2 Comparison of state variables from principal components and autoencoder
Figure C.7: Out-of-sample $R^{2}$ of return predictions and consensus approximations after compressing state variables to 32 dimensions from principal component analysis (PCA)
We next examine how the choice of macroeconomic feature compression method affects predictability. In particular, we compare autoencoder based compression with principal component analysis (PCA), both reduced to 32 dimensions. As shown in Figure C.7, PCA-based compression produces a pronounced decline in return predictability as $\lambda$ increases. This stands in sharp contrast to the autoencoder based compression, for which out-of-sample performance peaks around $\lambda=0.4$ and remains substantially higher overall. However, the two approaches deliver broadly similar performance for the consensus variable approximation. Taken together, these results indicate that, in our setting, the autoencoder provides a more effective representation of high-dimensional macroeconomic information for return prediction, even though both methods are comparable for consensus approximation. A plausible explanation is that PCA, being a linear and variance-based method, treats all input variables symmetrically and focuses solely on capturing overall variance, whereas the autoencoder can learn nonlinear transformations that emphasize features most relevant for the prediction task, thereby yielding more informative embeddings for returns.
C.4.3 Portfolio turnover and real-world implementability

### v2

@@ 6 Conclusion
A Consensus-Bottleneck Asset Pricing Model”  [^1] This paper is a revised version of Master’s thesis by Changeun Kim titled “A Consensus-Bottleneck Asset Pricing Model”, submitted to the Department of Industrial and Management Engineering, POSTECH, Korea. We would like to thank Hyeng Keun Koo (Committee), Kwangmin Jung (Committee), Dojoon Park (discussant), JinGi Ha, Jeonggyu Huh, Kyoung-Kuk Kim (discussant), Thummim Cho, and seminar participants at the 2024 Spring Joint Conference of Korean Operations Research and Management Science Society and Korean Institute of Industrial Engineers, 2025 Asia-Pacific Association of Finance International Conference, 2025 Korean Finance Association Fall Conference, 2025 4th Workshop on Financial Mathematics and Engineering (Pusan National University), 2025 Korea Derivatives Association Fall Conference, 2025 Annual Conference on Asia-Pacific Financial Markets (CAFM), for helpful discussions and insightful comments. This work was supported by the Ministry of Education of the Republic of Korea and the National Research Foundation of Korea [NRF-2023R1A2C2003927]. Jang (E-mail: bonggyujang@postech.ac.kr) is at Department of Industrial and Management Engineering, POSTECH, Korea University Business School; Jeong (E-mail: younwoo48@postech.ac.kr) is at Graduate School of Artificial Intelligence, POSTECH; Kim (E-mail: changeun120@postech.ac.kr) is at Department of Industrial and Management Engineering, POSTECH. Correspondence concerning this article should be addressed to Changeun Kim, Department of Industrial and Management Engineering, POSTECH, Pohang 37673, Republic of Korea. E-mail: changeun120@postech.ac.kr.
@@
The Internet Appendix is organized as follows. Section A reviews interpretable machine learning and situates CB-APM within this framework. Section B discusses data preprocessing procedures applied prior to estimation and evaluation. Section C provides details on the implementation and architectural choices of the neural network used throughout this study. Section D reports additional empirical results for robustness and supplementary insights. Section E presents the list of variables comprising the dataset used in this study.
Appendix A Interpretable Artificial Intelligence
Nevertheless the black box nature of deep neural networks, researchers may attempt to make such models understandable by introducing the concept of “disentangled representations” (See higgins2018towards; locatello2019challenging; locatello2019fairness). To understand the concept of disentanglement, we can consider the simple example of a naïve feedforward neural network, which consists of multiple hidden layers. The hidden layers are intermediate vector representations, often interpreted as features extracted from input data. Since they are trained with a downstream task-related objective in most cases, those layers are presumed to represent the essential information within the high-dimensional and noisy input data, well enough for the successful performance on that task. However, these features exist in an “entangled” manner, meaning that each element of the feature representation is a mixture of multiple factors. This entanglement complicates the interpretability of the model, as it blurs the specific contributions of individual input features to the final output. In this context, “disentanglement” refers to separating the underlying causal factors of the data and intermediate representations into distinct and non-overlapping representations. For instance, in fixed income markets, dozens of yields across maturities can be effectively summarized by three disentangled factors, level, slope, and curvature, as illustrated in Figure A.1. Each of these dimensions captures an independent source of variation in the yield curve, and together they provide an interpretable low-dimensional representation that still preserves the essential structure of the data. [^15] A clarification is needed to avoid conflating our framework with traditional factor models. nelson1987parsimonious impose level, slope, and curvature terms to parameterize the yield curve, while litterman1991common show via principal components that similar dimensions emerge empirically. Both approaches reduce high-dimensional yields to a few interpretable coordinates. By contrast, in CB-APM the consensus-bottleneck is not imposed ex ante but learned endogenously, closer in spirit to Litterman–Scheinkman than to Nelson–Siegel. In this sense, whereas Nelson–Siegel estimate factors to fit the curve, our model embeds interpretability within the network architecture itself, rendering part of the neural network transparent while retaining predictive capacity.
Figure A.1: Disentangled representations of neural network.
Appendix B Data Preprocessing
@@ B.3 Data imputation
However, for variables that represent growth rates or changes in firm characteristics,  [^17] Variables marked with asterisks in Table LABEL:tab:firm-level-predictors of Internet Appendix E we apply time-series mean imputation. This method captures the inherent continuity and trend in firm-specific dynamics, producing more realistic estimates than cross-sectional averaging. Relying repeatedly on the last observed value for growth-related factors could falsely imply persistence or monotonic trends, misrepresenting the inherently dynamic nature of such variables.
@@ B.4 Data normalization
Appendix C Implementing Neural Network for Asset Pricing
C.1 Activation functions
Rectified Linear Unit (ReLU) is frequently chosen in various machine learning applications due to its computational simplicity and efficiency, as also evidenced by its usage in empirical asset pricing research employing deep learning architectures such as gu2020empirical and chen2024deep. In contrast to activations such as SoftMax or Sigmoid functions, ReLU has demonstrated comparable performance while offering faster computational speed. Moreover, ReLU is particularly valued for its ability to address the gradient vanishing problem by consistently producing non-negative gradients for positive inputs. The Figure C.1 illustrates a graphical form of ReLU function.
Figure C.1: ReLU activation function.
@@ C.1 Activation functions
C.2 Stabilized learning
Stabilizing neural network training is essential for robust estimation, particularly when deploying deep learning models in high-stakes decision-making contexts. The following sections detail the regularization methods and machine learning techniques we employ to enhance the stability of the learning process. [^18] Detailed parameter settings are provided in Table LABEL:tab:hyper-setting. Despite that this topic is related to experimental factors rather than based on theoretical backgrounds, it is still worthy to discuss since that not only providing this information is crucial for reproducing the results, but also ensures a clear understanding of deep learning techniques. This is particularly important in fields like financial economics, where such concepts may not yet be widely understood or adopted. As such, this discussion is not only relevant but vital for integrating advanced computational methods into financial economic research effectively.
@@ C.2 Stabilized learning
Adaptive Moment Estimation (ADAM) is an optimization algorithm proposed by kingma2017adam, which is commonly used in practice of training deep learning models. It combines ideas from both momentum optimization and RMSprop (hinton2012neural), making it well-suited for optimizing non-convex objective functions commonly encountered in neural network training. ADAM maintains the first moment and the second moment of the gradients as two separate moving averages. These moving averages are used to adaptively update the parameters of the model during training. ADAM automatically adjusts the learning rate for each parameter based on the magnitude of the gradients and the accumulated past gradients, allowing it to converge quickly and efficiently in practice.
Learning rate scheduling is a technique used to dynamically adjust the learning rate during training to improve optimization performance. Instead of using a fixed learning rate throughout the training process, learning rate scheduling gradually decreases the learning rate over time, allowing the model to fine-tune its parameters more effectively as training progresses. By annealing the learning rate, learning rate scheduling helps to prevent the optimization process from getting stuck in local minima, which happens surprisingly often as discussed in the next section. Although ADAM inherently adjusts the learning rate, combining it with learning rate scheduling further optimizes performance by refining the initial learning rate. For training CB-APM, we use the ReduceLROnPlateau scheduler provided in PyTorch (See the official document provided by PyTorch for more details ReduceLROnPlateau), which monitors validation performance and reduces the learning rate by a specified factor.
@@ C.3 Model hyperparameters
Given these considerations, designing an appropriate hyperparameter optimization problem for asset pricing is important, which we leave it as an attractive and also challenging future research topic. In this paper, we rather summarize the hyperparameter “choice” that produces reasonable results as practical guidelines for finding a rational setting. The list of all the hyperparameters under consideration is provided in Table LABEL:tab:hyper-setting.
Table C.1: Hyperparameter settings for CB-APM and Autoencoder.
@@
Table C.1: Hyperparameter settings for CB-APM and Autoencoder (cont’d).
@@
Appendix D Additional Results
D.1 Forecast horizon analysis
Table D.1: Out-of-Sample $R^{2}$ for Stock Return and Consensus Approximations.
@@ D.1 Forecast horizon analysis
Table D.2: Out-of-Sample $R^{2}$ for Stock Return and Consensus Approximations.
@@
Table D.3: Out-of-Sample $R^{2}$ for Stock Return and Consensus Approximations.
@@
Figure D.1: Out-of-sample $R^{2}$ of return predictions and consensus approximations.
Panels (a), (b), and (c) present results for the monthly, quarterly, and semiannual returns respectively.
@@
Tables D.1D.3 and Figures 1(a)1(c) present the out-of-sample $R^{2}$ results and their period-by-period decomposition for these horizons. For shorter horizons such as one month, return predictability remains marginal, with $R^{2}$ values close to zero and even slightly negative at higher values of $\lambda$, consistent with the well-documented difficulty of forecasting near-term returns. In this regime, increasing $\lambda$ intensifies the interpretability-accuracy trade-off: while consensus approximation improves monotonically, return $R^{2}$ declines, suggesting that allocating more weight to consensus modeling diverts representational capacity away from short-horizon return-specific signals.
@@
D.2 Properties of the joint optimization
@@
Figure D.2 illustrates the in-sample MSE dynamics of the CB-APM under varying values of the hyperparameter $\lambda$, separately for monthly and annual forecasting horizons. Each panel presents two curves: the left axis depicts the in-sample MSE of stock return predictions, while the right axis reports the average in-sample MSE for consensus variable approximation. The first panel corresponds to the monthly horizon (Figure 2(a)), and the second panel presents the results for the annual horizon (Figure 2(b)).
@@
Figure D.2: In-sample MSE of return and consensus approximations.
This figure plots the in-sample MSE of stock return (left) and consensus approximation (right) for different $\lambda$ settings. Panel (a) reports results for the monthly return, and Panel (b) for the annual return. These plots illustrate how forecasting horizons and $\lambda$ values govern the trade-off between predictive accuracy and consensus reconstruction in the joint loss function.
The observed patterns reveal a striking divergence between the two horizons. For monthly returns, the stock return MSE exhibits a U-shaped trajectory: it initially decreases slightly for small values of $\lambda$ but increases steadily thereafter, suggesting a trade-off between return prediction accuracy and consensus approximation performance. This pattern aligns with the theoretical role of $\lambda$ in the loss function in equation (4), which explicitly prioritizes consensus approximation as its value increases. Placing greater weight on $\mathcal{L}_{C}$ directs more representational capacity of the network toward modeling analyst consensus at the expense of direct return prediction. This is consistent with the interpretability-accuracy trade-off widely documented in the interpretable machine learning literature (e.g., rudin2019stop), where models constrained to capture auxiliary structure or explanatory variables tend to sacrifice marginal predictive performance in favor of enhanced interpretability or alignment with economic reasoning.
@@
D.3 Structure of the learned macroeconomic representations
@@
Figure D.3 illustrates the two-dimensional principal component projection of the 32-dimensional latent state vectors produced by the macroeconomic autoencoder, color-coded by month and annotated with January observations for selected years. This visualization highlights how the autoencoder successfully encodes macroeconomic conditions into a smooth, low-dimensional manifold that evolves coherently over time. The trajectory of the latent vectors follows a clear temporal progression, demonstrating that the learned embedding captures the gradual transitions and structural shifts in the U.S. macroeconomic environment across the sample period.
Figure D.3: PCA projection of autoencoder latent state variables.
@@
A notable feature of this representation is its ability to distinguish major economic regimes. The grey dashed ovals in Figure D.3 correspond to periods classified as recessions by the National Bureau of Economic Research (NBER); the early 2000s recession (2001Q1–Q4) and the Global Financial Crisis (2007Q4–2009Q2). During these intervals, the latent vectors exhibit marked departures from their preceding trajectories, forming clusters that are distinct from surrounding expansionary phases. This pattern indicates that the autoencoder embedding effectively internalizes macroeconomic shocks and regime shifts, producing representations that align with well-established business cycle chronologies without direct supervision from recession labels.
@@
Recent deep factor models make it clear that neural networks can extract a parsimonious set of latent factors from high‐dimensional financial and macroeconomic data; these latent variables then drive improved asset pricing and predictive performance (see, for example, feng2018deep; gu2021autoencoder; chen2024deep). Our macroeconomic autoencoder performs a closely related function for aggregate time‐series data: it distills hundreds of macro indicators into a smooth latent trajectory that aligns with well‐known business‐cycle chronologies and regime shifts. While many prior studies emphasize quantitative performance metrics and offer only limited visual exploration of their latent factors, the clear temporal patterns in Figure D.3 demonstrate that CB‐APM’s autoencoder uncovers economically meaningful state dynamics from complex data.
@@
D.4 Further robustness checks
D.4.1 Sensitivity of autoencoder performance to latent dimensionality
@@
Table D.4: Out-of-Sample $R^{2}$ for Stock Return and Consensus Approximations for Different Autoencoder Dimensions.
@@
D.4.2 Comparison of state variables from principal components and autoencoder
Figure D.4: Out-of-sample $R^{2}$ of return predictions and consensus approximations after compressing state variables to 32 dimensions from PCA
We next examine how the choice of macroeconomic feature compression method affects predictability. In particular, we compare autoencoder based compression with PCA, both reduced to 32 dimensions. As shown in Figure D.4, PCA-based compression produces a pronounced decline in return predictability as $\lambda$ increases. This stands in sharp contrast to the autoencoder based compression, for which out-of-sample performance peaks around $\lambda=0.4$ and remains substantially higher overall. However, the two approaches deliver broadly similar performance for the consensus variable approximation. Taken together, these results indicate that, in our setting, the autoencoder provides a more effective representation of high-dimensional macroeconomic information for return prediction, even though both methods are comparable for consensus approximation. A plausible explanation is that PCA, being a linear and variance-based method, treats all input variables symmetrically and focuses solely on capturing overall variance, whereas the autoencoder can learn nonlinear transformations that emphasize features most relevant for the prediction task, thereby yielding more informative embeddings for returns.
D.4.3 Portfolio turnover and real-world implementability